A car accelerates from 4m/s to 20m/s in 8s. how far does it travel in time?


1
Expert's answer
2022-06-01T12:49:35-0400
"d=\frac{v_2+v_1}{2}t=\frac{20+4}{2}*8=96\:\rm m"